Nadal Company has 20 executives to whom it grants compensatory share options on January 1, 2019. At that time, it grants each executive the right to purchase 120 shares of its $5 par common stock at $40 per share after a 3-year service period. The value of each option is estimated to be $8.50 on the grant date. Based on its average employee turnover rate each year, Nadal expects that 2 executives will not vest in the plan. At the end of 2021, Nadal confirms that the actual turnover was the same as expected. On January 5, 2022, 5 executives exercise their options.

Required:

Prepare Nadal’s journal entries for 2019 through 2022 in regard to its compensatory share option plan.
Answer #1

Total estimated value of the options per employee = 120 * 8.50= $1020

Total extimated value of options for compensation expenses = $1020* 18 = $18360

Per year compensation expenses = $6120

Journal entries :

Date Accounts Titles Debit $

Credit$

Jan 1, 19 No entry
Dec 31, 19 Compensation Expenses 6120
Paid in Capital , stock options 6120
(18 * 120 * $8.50) / 3 = $6120
Dec 31, 20 Compensation Expenses 6120
Paid in Capital , stock options 6120
Dec 31, 21 Compensation Expenses 6120
Paid in Capital , stock options 6120
Jan 5, 22 Cash 11340
Paid in capital, stock options 3060
Common Stock 1800
Paid in Capital in excess of par, common stock 12600
(3 employees exercise their options)